The product of two consecutive even integers is 6 more than three times their sum .find the integers

Answer by stanbon(75887)

The product of two consecutive even integers
is 6 more than three times their sum.
find the integers
============================
1st: 2x
2nd: 2x+2
-----
Equation:
(2x)(2x+2) = 3(2x + 2x+2) + 6
4x^2+4x = 3(4x+2) + 6
4x^2+4x = 12x + 12
4x^2 - 8x -12 = 0
---
x^2 - 2x -3 = 3
-----
(x-3)(x+1) = 0
x = 3 or x = -1
----
Positive solution:
1st 2*3 = 6
2nd: 2x+2 = 8
===================
Cheers,
Stan H.
===================